# Torsion and Chern-Simons gravity in 4D space-times from a Geometrodynamical four-form

### v2

@@ Abstract
The particle aspects of the geometrodynamical four-form field is considered and is shown to be a pseudo-scalar. It is shown that the Bose-Einstein condensates of such pseudo-scalars can give rise to formation of supermassive black holes through an interplay of self-gravity and quantum mechanics. Therefore, self-gravitating BECs of pseudo-scalar particles associated with the geometrodynamical four-form field may solve the long standing problem concerning frequent discovery of tens of billion solar mass SMBHs at epochs when the universe was barely a billion years old.
It is also demonstrated that in the presence of the dynamical four-form, gravitational waves can be decoupled using circularly polarized waveforms. Moreover, exact, self-consistent monochromatic gravitational wave solutions of the wave equation can be obtained. However, much more work is needed to obtain realistic solutions of gravitational waves, with group velocities $\leq c$, propagating in the background consisting of $\tilde{w}$.
VII Acknowledgements
Acknowledgements.

### v3

@@ 6 Conclusions
The geometrodynamical four-form field is shown to correspond to a pseudo-scalar particle. From a simplified semi-classical treatment, it is demonstrated that the Bose-Einstein condensates of such pseudo-scalars can give rise to formation of supermassive black holes through an interplay of self-gravity and quantum mechanics. Therefore, self-gravitating BECs of pseudo-scalar particles associated with the geometrodynamical four-form field may solve the long standing problem concerning the frequent discoveries of tens of billion solar mass SMBHs at epochs when the universe was less than billion years old.
It is also proved that in the presence of the dynamical four-form, equations for the linearly polarized gravitational waves can be decoupled using circularly polarized waveforms as well as exact and self-consistent solutions of the dynamical four-form equation coupled to a monochromatic gravitational wave (albeit with phase velocity $>c$) can be obtained. However, much more work is needed to obtain realistic solutions of gravitational waves, with group velocities $\leq c$, propagating in the background consisting of $\tilde{w}$.
7 Acknowledgements